the equation for a projectile height versus time is h(t)= -16^2+ Vot+ho

a tennis ball machine serves a ball vertically into the air from a height of two feet, with an initial speed of 120 feet per second . after how many seconds does the ball attain its maximum height


1
Expert's answer
2022-03-02T12:13:27-0500
"h(t)=2+120t-16t^2, t\ge0"

Find the first derivative with respect to "t"


"h'(t)=120-32t"

Find the critical number(s)


"h'(t)=0=>120-32t=0"

"t=3.75\ s"


Critical number: 3.75


If "0\le t<3.75, h'(t)>0, h(t)" increases.

If "t>3.75, h'(t)<0, h(t)" decreases.

The function "h(t)" has a local maximum at "t=3.75." Since the function "h(t)" has the only extremum for "t\ge0," then the function "h(t)" has the absolute maximum at "t=3.75."

The ball attains its maximum height after "3.75" seconds .